OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the following point: y=(1+6x)^10 (0,1)

OpenStudy (anonymous):

do i have to use the equation y-y1=m(x-x1)

OpenStudy (anonymous):

@91004775 yes, woops, I thought this problem was trivial.$$y=(1+6x)^{10}\\y'=60(1+6x)^9\\y'(0)=60$$So we have the point \((0,1)\) and the slope \(60\), meaning:$$\frac{y-1}{x-0}=60\\y-1=60(x-0)=60x\\y=60x+1$$

OpenStudy (anonymous):

okay yeah that's what i got thank you :)